### v1

@@ 2.1 Problem formulation
The interior of the entire computational domain is a bounded, open, and connected region $\Omega\subset\mathbb{R}^{d}$, and $\Omega\subset\mathbb{R}^{d}$ has a polygonal boundary $\partial\Omega$. The system is governed by the operator $\mathcal{A}$ and a source term $f(\bm{x})$ within the domain $\Omega$, subject to boundary conditions on $\partial\Omega$ prescribed by the operator $\mathcal{B}$ and the function $g(\bm{x})$.
@@ 2.2 Introduction to PINNs
where $V(\Omega)$ is the volume of $\Omega$. Without loss of generality, we use $\alpha_{1}$ to denote the weight of the discrete residual loss function $\mathcal{L}_{r,N_{r}}(\bm{x}_{i};\theta)$, and $\alpha_{2}$ to denote the weight of the discrete boundary loss function $\mathcal{L}_{b,N_{b}}(\bm{x}_{i};\theta)$. Ultimately, the optimization method is employed to minimize the empirical loss $\mathcal{L}_{N}(\bm{x};\theta)$, steering PINNs toward learning the latent solution of the equation 1.
@@ 3.1.1 Multi-level Sampling
In this section, we will elaborate on the sampling algorithm within the Multi-Level framework, which named as Multi-Level sampling. The vision of this sampling method is that sampling points can be concentrated in areas with large loss residuals and drastic changes in the predicted solution, and become increasingly concentrated as the level increases. This indicates that the Multi-Level sampling method will be used multiple times in our framework. Below, we will start with the introduction of this algorithm from the first-level.
@@
$\mathcal{H}(\bm{x};\theta^{1})=\gamma_{1}|\nabla u(\bm{x};\theta^{1})|,$ (5)
where $\gamma_{1}$ is a given constant. Following the guidance of Eq (4), sampling points can be concentrated in regions with larger residuals and sharp gradients. Similarly, Eq (5) enables the concentration of sampling points according to the properties of the predicted solution $u(\bm{x};\theta^{1})$. Therefore, we construct the following monitor function.
$\mathcal{W}(\bm{x};\theta^{1})=\frac{\left(\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right)^{\frac{1}{2}}}{\int_{\Omega}\left(\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right)^{\frac{1}{2}}d\bm{x}}.$ (6)
Finally, the second-level residual training points $\mathcal{X}^{2}_{N_{r}}=\left\{x_{i}\right\}_{i=1}^{N_{r}}$ can be sampled according to the density function $p(\bm{x};\theta^{1})$ of the monitoring function $\mathcal{W}(\bm{x};\theta^{1})$.
$p(\bm{x};\theta^{1})\propto\mathcal{W}(\bm{x};\theta^{1})+1.$ (7)
@@
$\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)=\gamma_{1}|\nabla[u(\bm{x};\theta^{1})+u(\bm{x};\theta^{2})]|.$ (9)
@@
$\mathcal{W}\left(\bm{x};(\theta^{1},\theta^{2})\right)=\mu\mathcal{W}\left(\bm{x};\theta^{1}\right)+\frac{\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{\frac{1}{2}}}{\int_{\Omega}\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{\frac{1}{2}}d\bm{x}},$ (10)
where $\mu$ is a give postive constant. Thus, the corresponding density function is obtained from the updated monitoring function to sample the third - level training points $\mathcal{X}^{3}_{N_{r}}=\left\{x_{i}\right\}_{i=1}^{N_{r}}$.
@@
The details of the subsequent-level sampling will not be introduced one by one but will be summarized in Algorithm 1.
@@ 1. Update the residual function:
$\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)$ $=\gamma_{1}|\nabla[u(\bm{x};\theta^{1})+u(\bm{x};\theta^{2})+.+u(\bm{x};\theta^{k}]|$
@@
$\mathcal{W}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)$ $=(k-1)\mu\mathcal{W}\left(\bm{x};\theta^{1}\right)+(k-2)\mu\mathcal{W}\left(\bm{x};(\theta^{1},\theta^{2})\right)$
@@ 4.1 Symbols and parameter settings
In order to reduce the effect of randomness on the calculation results, we set all the seeds of the random numbers to 100. The default parameter settings in PINNs are given in Table 1. All numerical simulations are carried on NVIDIA A100 with 80GB of memory and NVIDIA A800 with 80GB of memory.
@@ 4.2 Two-dimensional Poisson equation
In this experiment, we sample 10000 points within $\Omega$ as the residual training set and 1000 points on $\partial\Omega$ as the boundary training set, while utilizing $400\times 400$ points for the test set, and we employ a four-level framework for training. In the first-level pre-training, we trained 20000 epochs using the SOAP method and 10000 epochs using the SSB method. In the second-level training, we again trained 20000 epochs of the SOAP method and 10000 epochs of the SSB method. In the third-level training, we employed 5000 epochs of the SSB method. Finally, in the fourth-level training, we used 10000 epochs of the SSB method. During training at the different levels, we employed the MLS method described in Section 3.1.1 and the MLT method described in Section 3.1.2.

### v2

@@ 2.1 Problem formulation
The interior of the entire computational domain $\Omega\subset\mathbb{R}^{d}$ is a bounded, open, and connected region , and it has a polygonal boundary $\partial\Omega$. The system is governed by the operator $\mathcal{A}$ and a source term $f(\bm{x})$ within the domain $\Omega$, subject to boundary conditions on $\partial\Omega$ prescribed by the operator $\mathcal{B}$ and the function $g(\bm{x})$.
@@ 2.2 Introduction to PINNs
where $V(\Omega)$ is the volume of $\Omega$. Without loss of generality, we use $\alpha_{1}$ to denote the weight of the discrete residual loss function $\mathcal{L}_{r,N_{r}}(\bm{x}_{i};\theta)$, and $\alpha_{2}$ to denote the weight of the discrete boundary loss function $\mathcal{L}_{b,N_{b}}(\bm{x}_{i};\theta)$. Ultimately, the optimization method is employed to minimize the empirical loss $\mathcal{L}_{N}(\bm{x};\theta)$, steering PINNs toward learning the latent solution of the Eq (1).
@@ 3.1.1 Multi-level Sampling
In this section, we will elaborate on the sampling algorithm within our Multi-Level depp framework, which named as Multi-Level sampling. The vision of this sampling method is that sampling points can be concentrated in areas with large loss residuals and drastic changes in the predicted solution, and become increasingly concentrated as the level increases. This indicates that the Multi-Level sampling method will be used multiple times in our framework. Below, we will start with the introduction of this algorithm from the first-level.
@@
$\mathcal{H}(\bm{x};\theta^{1})=\gamma|\nabla u(\bm{x};\theta^{1})|,$ (5)
where $\gamma$ is a given constant. Following the guidance of Eq (4), sampling points can be concentrated in regions with larger residuals and sharp gradients. Similarly, Eq (5) enables the concentration of sampling points according to the properties of the predicted solution $u(\bm{x};\theta^{1})$. Therefore, we construct the following monitor function.
$\mathcal{W}(\bm{x};\theta^{1})=\frac{\left(\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right)^{s}}{\int_{\Omega}\left(\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right)^{s}d\bm{x}},$ (6)
where the exponent s is a positive parameter. A larger value of s leads to a higher concentration trend. Finally, the second-level residual training points $\mathcal{X}^{2}_{N_{r}}=\left\{x_{i}\right\}_{i=1}^{N_{r}}$ can be sampled according to the density function $p(\bm{x};\theta^{1})$ of the monitoring function $\mathcal{W}(\bm{x};\theta^{1})$.
$p(\bm{x};\theta^{1})\propto\mathcal{W}(\bm{x};\theta^{1})+c.$ (7)
Where c is a penalty term constant, which ensures that even in regions where the monitoring function is small, a sufficient number of sampling points will still be obtained during sampling.
@@
$\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)=\gamma|\nabla[u(\bm{x};\theta^{1})+u(\bm{x};\theta^{2})]|.$ (9)
@@
$\mathcal{W}\left(\bm{x};(\theta^{1},\theta^{2})\right)$ $=\mu\frac{\left[\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right]^{s}}{\int_{\Omega}\left[\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right]^{s}d\bm{x}}+\frac{\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{s}}{\int_{\Omega}\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{s}d\bm{x}}$ (10)
$=\mu\mathcal{W}(\bm{x};\theta^{1})+\frac{\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{s}}{\int_{\Omega}\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{s}d\bm{x}},$
where $\mu$ is a given postive constant. The presence of $\mu$ ensures that as the level increases, the points become more concentrated. Thus, the corresponding density function is obtained from the updated monitoring function to sample the third - level training points $\mathcal{X}^{3}_{N_{r}}=\left\{x_{i}\right\}_{i=1}^{N_{r}}$.
@@
The details of the subsequent-level sampling will not be introduced one by one but will be summarized in Algorithm 1. For details on the parameter settings in the MLS method, please refer to Table 1.
@@ 1. Update the residual function:
$\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)$ $=\gamma|\nabla[u(\bm{x};\theta^{1})+u(\bm{x};\theta^{2})+.+u(\bm{x};\theta^{k}]|$
@@
$\mathcal{W}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)$ $=(k-1)\mu\frac{\left[\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right]^{s}}{\int_{\Omega}\left[\mathcal{R}(\bm{x};\theta^{1})+\mathcal{H}(\bm{x};\theta^{1})\right]^{s}d\bm{x}}+(k-2)\mu\frac{\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{s}}{\int_{\Omega}\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2})\right)\right]^{s}d\bm{x}}$
@@
$=\mu\mathcal{W}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k-1})\right)+\frac{\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)\right]^{\frac{1}{2}}}{\int_{\Omega}\left[\mathcal{R}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)+\mathcal{H}\left(\bm{x};(\theta^{1},\theta^{2},...,\theta^{k})\right)\right]^{\frac{1}{2}}d\bm{x}}$
@@ 3.2 Error Estimation
Assumption 2.
Let $\forall\bm{u}\in\mathbb{H}$, the following assumption holds
$C_{1}\|\bm{u}\|_{2}\leq\|\mathcal{A}(\bm{u})\|_{2}+\|\mathcal{B}(\bm{u})\|_{2}\leq C_{2}\|\bm{u}\|_{2},$ (35)
where $C_{1}$ and $C_{2}$ are positive constants.
In order to give an upper bound of the error, we need to introduce the the Rademacher complexity.
Definition 3.
Let $\mathcal{X}:=\{x_{i}\}_{i=1}^{n}$ be a collection of i.i.d. random samples, the Rademacher complexity of the function class $\mathcal{F}$ can be defined as
$\mathcal{R}_{n}(\mathcal{F})=\mathbb{E}_{\mathcal{X},\epsilon}\left(\sup_{f\in\mathcal{F}}|\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(x_{i})|\right),$
where $(\epsilon_{1},\epsilon_{2},...,\epsilon_{n})$ are i.i.d. Rademacher variables, i.e., $P(\epsilon_{i}=1)=P(\epsilon_{i}=-1)=0.5$.
Now we consider the approximate solution in the set $\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1})=\left\{\bm{u}_{k}+\sum_{i=1}^{k-1}\bm{u}_{i}|\forall\bm{u}_{k}\in D_{1}^{NN}\right\}$.
Assumption 3.
$\forall\bm{u}\in\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1})$, there exists a positive number $0<B_{r}\in\mathbb{R}$ such that
$\sup\|\mathcal{A}[\bm{u}(\bm{x})]-f(x)\|_{\infty}=B_{r},$ (36)
$\sup\|\mathcal{B}[\bm{u}(\bm{x})]-g(x)\|_{\infty}=B_{g}.$
According to [20], we use Lemma 2 and propose the following theorem about the Rademacher complexity.
Lemma 2.
Let $\mathcal{F}$ be a b-uniformly bounded class of integrable real-valued functions with domain $\Omega$, for any postive integer $n\geq 1$ and any scalar $\delta\geq 0$, we have
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}f(x_{i})-\mathbb{E}[f(x)]\right|\leq 2\mathcal{R}_{n}(\mathcal{F})+\delta,$
with probability at least $1-exp\left(-\frac{n\delta^{2}}{2b^{2}}\right)$.
Theorem 3.
Suppose that Assumption 2 and Assumption 3 hold, let $\bm{u}^{*}$ be the exact solution of Eq (1). $\forall\delta>0$, $\{x_{i}\}_{i=1}^{M_{r}}$ is a collection of i.i.d. random samples and $\forall\bm{u}(\bm{x};\theta)\in\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1})$, the following error estimate holds
$\|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)\|_{2}\leq\frac{\sqrt{2}}{C_{1}}\left(\|r(\bm{x};\theta)\|_{2,N_{r}}^{2}+\|b(\bm{x};\theta)\|_{2,N_{b}}^{2}+2\mathcal{R}_{N_{r}}(\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1}))+2\mathcal{R}_{N_{b}}(\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1}))+\delta\right)^{\frac{1}{2}},$ (37)
with probability at least $\left(1-\exp\left(-\frac{N_{r}\delta^{2}}{8B_{r}^{2}}\right)\right)\left(1-\exp\left(-\frac{N_{b}\delta^{2}}{8B_{g}^{2}}\right)\right)$.
Proof.
By Assumption 2, we have
$C_{1}\|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)\|_{2}$ $\leq\|\mathcal{A}(\bm{u}^{*}(\bm{x}))-\mathcal{A}(\bm{u}(\bm{x};\theta))\|_{2}+\|\mathcal{B}(\bm{u}^{*}(\bm{x}))-\mathcal{B}(\bm{u}(\bm{x};\theta))\|_{2}$ (38)
$=\|\mathcal{A}(\bm{u}(\bm{x};\theta))-f\|_{2}+\|\mathcal{B}(\bm{u}(\bm{x};\theta))-g\|_{2}$
$=\|r(\bm{x};\theta)\|_{2}+\|b(\bm{x};\theta)\|_{2}$
$\leq\sqrt{2}\left(\|r(\bm{x};\theta)\|_{2}^{2}+\|b(\bm{x};\theta)\|_{2}^{2}\right)^{\frac{1}{2}}.$
Using Lemma 2, we have
$\|r(\bm{x};\theta)\|_{2}^{2}\leq\|r(\bm{x};\theta)\|_{2,N_{r}}^{2}+2\mathcal{R}_{N_{r}}(\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1}))+\frac{\delta}{2},$ (39)
with probability at least $1-\exp\left(-\frac{N_{r}\delta^{2}}{8B_{r}^{2}}\right)$.
$\|b(\bm{x};\theta)\|_{2}^{2}\leq\|b(\bm{x};\theta)\|_{2,N_{b}}^{2}+2\mathcal{R}_{N_{b}}(\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1}))+\frac{\delta}{2},$ (40)
with probability at least $1-\exp\left(-\frac{N_{b}\delta^{2}}{8B_{g}^{2}}\right)$.
Therefore,
$\|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)\|_{2}$ $\leq\frac{\sqrt{2}}{C_{1}}(\|r(\bm{x};\theta)\|_{2}^{2}+\|b(\bm{x};\theta)\|_{2}^{2})^{\frac{1}{2}}$ (41)
$\leq\frac{\sqrt{2}}{C_{1}}\left(\|r(\bm{x};\theta)\|_{2,N_{r}}^{2}+\|b(\bm{x};\theta)\|_{2,N_{b}}^{2}+2\mathcal{R}_{N_{r}}(\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1}))+2\mathcal{R}_{N_{b}}(\mathcal{F}^{NN}_{k}(\bm{u}_{1},...,\bm{u}_{k-1}))+\delta\right)^{\frac{1}{2}},$
with probability at least $\left(1-\exp\left(-\frac{N_{r}\delta^{2}}{8B_{r}^{2}}\right)\right)\left(1-\exp\left(-\frac{N_{b}\delta^{2}}{8B_{g}^{2}}\right)\right)$. ∎
@@ 4.1 Symbols and parameter settings
In order to reduce the effect of randomness on the calculation results, we set all the seeds of the random numbers to 100. The default parameter settings in MLS method and PINNs are given in Table (1 and 2). All numerical simulations are carried on NVIDIA A100 with 80GB of memory and NVIDIA A800 with 80GB of memory.
Table 1: Default settings for main parameters in MLS method.
    $\gamma$     s     c     $\mu$
    1     0.5     1     4
@@ 2.4.0 Tanh Xavier 1/1 $40\times 4$ 100
4.2 One-dimensional Advection-Diffusion equation
Considering the following one-dimensional convection-diffusion equation
[…]
where $\Omega=(-1,1)$. When $f(x)\equiv 1$, the solution to this problem is given by
$\hskip-8.5359pt\begin{array}[]{r@{}l}\begin{aligned} u=\frac{x}{b}+\frac{1}{b}\left(\frac{e^{-\frac{2b}{\epsilon}+1}}{e^{-\frac{2b}{\epsilon}-1}}\right)\left(\frac{2e^{\frac{b}{\epsilon}(x-1)}}{e^{-\frac{2b}{\epsilon}}+1}-1\right).\end{aligned}\end{array}$ (44)
When the value of $\epsilon$ is small, it is difficult to solve the boundary layer accurately. Therefore, we set $\epsilon=0.01$ and b=1b=-1.
In this experiment, we sample 200 points within $\Omega$ as the residual training set, and we employ a three-level framework for training. In the first-level pre-training, we trained 5000 epochs using the SOAP method. Then, in the second-level training, we again trained 5000 epochs of the SOAP method and 3000 epochs of the SSB method. Finally, in the third-level training, we employed 3300 epochs of the SSB method.
During training at the different levels, we employed the MLS method described in Section 3.1.1. Fig 3 illustrates the effect of the MLS method across the various levels; it can be seen that as the level increases, the points progressively concentrate toward x=1x=-1, because the solution has very low regularity in the vicinity of x=1x=-1. It can be seen that the MLS method is able to capture this property effectively.
(a) 1st level
(b) 2nd level
(c) 3rd level
Figure 3: The sampling points at different levels for the one-dimensional advection-diffusion equation with the solution Eq (44). This figure shows the behavior of sampling points in [1,0.5][-1,-0.5].
The numerical results of different levels are given in Fig 4. Fig 4 illustrates figures based on the predicted solutions $\bm{u}(\bm{x};\theta)|$ and the absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of our multi-level deep framework.
(a) 1st prediction
(b) 2nd prediction
(c) 3rd prediction
(d) 1st error
(e) 2nd error
(f) 3rd error
Figure 4: The numerical result of different levels for the one-dimensional advection-diffusion equation with the solution Eq (44). (a)–(c) Comparison between the predicted $\bm{u}(\bm{x};\theta)$ and exact solutions $\bm{u}^{*}(\bm{x})$. (d)–(f) Performance of the absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$.
In Fig 5, we compare the commonly used RAD method in PINNs under the same number of training epochs. Specifically, the training process consists of pre-training with 5000 iterations of Adam optimizer, followed by sampling with the RAD method, finally trains again for 5000 Adam epochs and 6300 L-BFGS epochs. Fig 5 illustrates the variation of the relative errors during the training process, from which the superiority of our method can be observed. In the subsequent experiments, we will conduct a more detailed comparison on more complex cases. Ultimately, the relative errors of our method in solving the one-dimensional advection-diffusion equation with the solution Eq (44) are $e_{\infty}(\bm{u})=1.561\times 10^{-7}$ and $e_{2}(\bm{u})=5.494\times 10^{-8}$.
(a) performance of $e_{\infty}(u)$
(b) performance of $e_{2}(u)$
Figure 5: The performance of errors for the one-dimensional advection-diffusion equation with the solution Eq (44). (a) the relative error $e_{\infty}(u)$ with different training epochs; (b) the relative error $e_{2}(u)$ with different training epochs.
@@ 4.3 Two-dimensional Poisson equation
In this experiment, we sample 10000 points within $\Omega$ as the residual training set and 1000 points on $\partial\Omega$ as the boundary training set, while utilizing $400\times 400$ points for the test set, and we employ a four-level framework for training. In the first-level pre-training, we trained 20000 epochs using the SOAP method and 10000 epochs using the SSB method. In the second-level training, we again trained 20000 epochs of the SOAP method and 10000 epochs of the SSB method. In the third-level training, we employed 5000 epochs of the SSB method. Finally, in the fourth-level training, we used 10000 epochs of the SSB method.
During training at the different levels, we employed the MLS method described in Section 3.1.1. Fig 7 illustrates the effect of the MLS method across the various levels; it can be seen that as the level increases, the points progressively concentrate toward the origin, because the solution has very low regularity in the vicinity of the origin. It can be seen that the MLS method is able to capture this property effectively.